john read 66 books in 12 months at this rate how many books will he read in 4 months
Found 2 solutions by stanbon, guitarheroguy:
Answer by stanbon(75887)

john read 66 books in 12 months at this rate how many books will he read in 4 months
x/4 = 66/12
----
x = 4(66/12)
---
x = 22
-------------------
Cheers,
Stan H.

Answer by guitarheroguy(26)

Well...
We need to set up a proportion. Here is how we do it.
4=12
x=66
Now we cross multiply.
4*66=264
12*x=12x
Now to isolate the variable by dividing both sides by 12
12x/12=x
264/12=22
x=22
John will read 22 books in 4 months